OpenStudy (anonymous):

if 3x=y-5 and x=-4 then y=?

OpenStudy (anonymous):

-12=y-5 -7=y

OpenStudy (anonymous):

3*(-4)=y-5

OpenStudy (petiteme):

Substitute the value -4 to your x